OpenStudy (anonymous):

Improper Integrals: Evaluate integral from 0 to pi/2 of tan θ dθ, or show that it is divergent.

OpenStudy (anonymous):

I've gotten an answer here, that is is divergent because sec (pi/2)^2 does not exist.

OpenStudy (zarkon):

more importantly \[\lim_{t\to\frac{\pi}{2}^{-}}\sec^2(t)=\infty\]

OpenStudy (anonymous):

Thanks.